OpenStudy (anonymous):

Solve for J: 2/5 + J = 4/5.

OpenStudy (gorv):

2/5

OpenStudy (anonymous):

J= \[\frac{ 4 }{5 } - \frac{ 2 }{ 5 }\] So J = \[\frac{ 2 }{ 5 }\]